You have prepared a solution by dissolving 14.659g of dodecane (C12H26) in 40.00mL of cyclohexane (C6H12) at 25°C. What is the expected vapor pressure (in units of mmHg) of this solution? At 25°C, the density of cyclohexane is 0.7792g/mLand its vapor pressure is 97.5mmHg. Assume that dodecane is a non-volatile solute.

Your Answer:

Question 11 options:

Answer units
Answer #1

Here is the solution of your question. If you have any doubt or need any clarification please comment in comment box and will definitely resolve your query. If you find useful please upvote it. Thanks in advance.

mole of dodecane = 14-659 gm 170gm/mol 0.086 mol 0.7792gm x 40ml = 31-17 gm mass of cycloberane = dXV = mole of 31-17gm Cyclo